OpenStudy (anonymous):

if (3x^2)+7=0 then (x-1/3)^2 how do you suppose to solve this problem?

OpenStudy (anonymous):

Well for the first equation we have 3x^2 + 7 = 0 so the the (3x^2) part needs to equal -7 to make the equation work. So we would do 3x^2 = 7 . Can you find x from that? And then you would just plug X into the next equation to find the answer.

OpenStudy (anonymous):

I made a mistake. 3x^2 = -7 not 7 my bad

OpenStudy (anonymous):

so the first expression is needed to solve for the second one?

OpenStudy (anonymous):

let me solve

OpenStudy (anonymous):

Yes, because you need to make it equal to 0? What class is this for? I am not sure how far you need to go with this. Cuase you would maybe have to use imaginary numbers using i

OpenStudy (anonymous):

it is a problem from a practice test i know the answer by looking it up on the answer key but I would like to know the steps

OpenStudy (anonymous):

And is the question states if 3x^2+7 = 0 then what is the value of ((x-1)/3)^2

OpenStudy (anonymous):

is that the question ^ because what u said is a little different

OpenStudy (anonymous):

if 3x^2-2x+7=0, then (x-1/3)^2

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

3x^2 + 7 = 0 3x^2 = -7 3x = square root of -7 which then you would have to use imaginary numbers 3x = isquareroot of 7 x = i square root of 7/3

OpenStudy (anonymous):

And i do not understand so the queston is if 3x^2-2x+7=0, then (x-1/3)^2?? but why are they giving you the second equation, do they want the value of (x-1/3)^2

OpenStudy (anonymous):

the answer is -20/9

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

I dont know sorry, the question just doesnt make sense, i dont know what they are looking for.

OpenStudy (anonymous):

ok, thanks for your help anyhow

OpenStudy (anonymous):

\[3x ^{2}-2x+7=0,3\left( x ^{2}-\frac{ 2 }{ 3 }x+\frac{ 1 }{ 9 }-\frac{ 1 }{9} \right)+7=0\] \[3\left( x-\frac{ 1 }{ 3 } \right)^{2}-3*\frac{ 1 }{9 }+7=0\] \[3\left( x-\frac{ 1 }{ 3 } \right)^{2}=\frac{ 1 }{3}-7=-\frac{ 20 }{3 }\] \[\left( x-\frac{ 1 }{ 3 } \right)^{2}=-\frac{ 20 }{9 }\]

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

if i can ask, -1/9 & 1/9 come from 1/3^2

OpenStudy (anonymous):

how come one is negative and one positive

OpenStudy (anonymous):

\[\to complete thesquare ,if coefficient of x ^{2} is 1,then add \left( \frac{coefficient of x}{ 2 } \right)^{2}\]

OpenStudy (anonymous):

ok!